# EW corrections and Heavy Boson Radiation at a high-energy muon collider

### v2

@@ 4.3.2 EW jets in multi-$V$ production
We now consider the case of inclusive EW-dijet production, $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, where in the EW-jet $j_{\rm EW}$ we cluster $W$ and $Z$ bosons. In other words, the LO originates from the processes $\mu^{+}\mu^{-}\to VV$ with $V=W,Z$, which correspond to the final states $WW$ and $ZZ$. Having found in the previous section that the impact of the Higgs boson radiation is minimal, we consider as a contribution to the HBR only the processes involving an additional $V=W,Z$ radiation, $\mu^{+}\mu^{-}\to 3V$, which correspond to the final states $ZZZ$ and $WWZ$. Clearly, the (at least) two $j_{\rm EW}$’s that are required may be also associated with $W$ and $Z$, respectively, or also with a $WZ$ pair and a $W$ respectively. Several combinations are possible and some of them, as the ones just mentioned and unlike the cases discussed in the previous section, do not feature the same “partonic” final state of the LO, which can feature only $ZZ$ and $WW$. We will consider also the case of a double HBR and therefore the $4V$ final states: $ZZZZ$, $WWWW$ and $WWZZ$ [^25] As a side comment, such processes have been shown to be very sensitive to possible anomalous interactions of the Higgs boson with the muon Celada:2023oji ; Han:2021lnp and therefore are of particular relevance in the muon collider physics program.
Figure 23: The $p_{T}(j_{{\rm EW},1})$ distribution in $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, with $m(2j_{\rm EW})>0.8\sqrt{S}$. The left (right) plot shows results at $\sqrt{S}=3$ TeV ($\sqrt{S}=10$ TeV). The histograms show LO (dashed) and NLO (solid predictions) for $W^{+}W^{-}$ (green), $ZZ$ (blue), $2V=W^{+}W^{-}+ZZ$ (orange), $3V$ (red), and $4V$ (violet, only at the LO). In the second inset, the quantities $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$ and $\delta_{\rm NLO_{EW}+HBR}$ are shown respectively as black-solid and black-dashed lines. These quantities are defined by Eq. (3.26) in terms of respectively Eq. (3.25) and Eq. (3.24). The solid grey line shows the quantity $\delta_{{\rm SDK_{weak}}}^{2}/2$, which enters only in $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$.
@@
In the main panel we notice that the LO is dominated by the $WW$ “partonic” process, which has a much larger cross section than the $ZZ$ “partonic” process and therefore coincides with $2V$. Especially at 10 TeV, both $2V$ and $3V$ processes receive very large and negative contributions from NLO EW corrections. The coloured dotted lines in the main inset of the right plot of Fig. 23 indicates negative values for the lines depicted as solid, which we plot in absolute value in order to be visible on a logarithmic scale. It means that for both $2V$ and $3V$ processes, the relative NLO EW corrections are negative and larger than $100\%$ in absolute value. At 3 TeV they are smaller, but also in this case they are much larger than the contribution from double HBR production at LO, i.e. $4V$ at LO. Similarly to the case of the HBR, i.e. $3V$ at LO, which is smaller than the NLO EW corrections, the double HBR, i.e. $4V$ at LO, is smaller than the NLO EW corrections to $3V$ and so to the single HBR. Estimating the two-loop corrections, i.e. the NNLO to $2V$, as $\delta_{\rm NLO_{EW}}^{2}/2\simeq\delta_{{\rm SDK_{weak}}}^{2}/2$, we see that this contribution (the grey solid line) is even larger than the NLO EW corrections to the HBR contribution. In other words, looking at the case ${\cal O}(\alpha)$ and ${\cal O}(\alpha^{2})$, it appears that for a given $n$ the leading component of the ${\cal O}(\alpha^{n})$ corrections to LO in $2j_{\rm EW}$ production is the $n$-loop corrections, the subleading component is the $(n-1)$-loop corrections to HBR and so on; loops win over HBR.
@@ Appendix A ISR parton distribution
consistently with the renormalisation scheme that we will employ in the computations, with $G_{\mu}$ measured from the muon decay. Finally, as it is well known, the usage of LL-accurate ISR in NLO EW computations requires the inclusion, at the level of short-distance cross section, of additional terms required to attain formal NLO accuracy (treated on the same footage as change of factorisation-scheme contributions) 272727 For more details, see e.g. Sec. 2 of Ref. Denner:2003iy or Appendix A of Ref. Bertone:2022ktl . . All our results at NLO will be computed including these terms, which are already available inside MadGraph5_aMC@NLO.

### v3

@@ 4.3.2 EW jets in multi-$V$ production
We now consider the case of inclusive EW-dijet production, $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, where in the EW-jet $j_{\rm EW}$ we cluster $W$ and ZZ bosons. In other words, the LO originates from the processes $\mu^{+}\mu^{-}\to VV$ with $V=W,Z$, which correspond to the final states $WW$ and $ZZ$. Having found in the previous section that the impact of the Higgs boson radiation is minimal, we consider as a contribution to the HBR only the processes involving an additional $V=W,Z$ radiation, $\mu^{+}\mu^{-}\to 3V$, which correspond to the final states $ZZZ$ and $WWZ$. Clearly, the (at least) two $j_{\rm EW}$’s that are required may be also associated with $W$ and $Z$, respectively, or also with a $WZ$ pair and a $W$ respectively. Several combinations are possible and some of them, as the ones just mentioned and unlike the cases discussed in the previous section, do not feature the same “partonic” final state of the LO, which can feature only $ZZ$ and $WW$. We will consider also the case of a double HBR and therefore the $4V$ final states: $ZZZZ$, WWWWWWWW and WWZZWWZZ. [^25] As a side comment, such processes have been shown to be very sensitive to possible anomalous interactions of the Higgs boson with the muon Celada:2023oji ; Han:2021lnp and therefore are of particular relevance in the muon collider physics program.
Figure 23: The $p_{T}(j_{{\rm EW},1})$ distribution in $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, with $m(2j_{\rm EW})>0.8\sqrt{S}$. The left (right) plot shows results at $\sqrt{S}=3$ TeV ($\sqrt{S}=10$ TeV). The histograms show LO (dashed) and NLO (solid predictions) for $W^{+}W^{-}$ (green), ZZZZ (blue), $2V=W^{+}W^{-}+ZZ$ (orange), $3V$ (red), and $4V$ (violet, only at the LO). In the second inset, the quantities $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$ and $\delta_{\rm NLO_{EW}+HBR}$ are shown respectively as black-solid and black-dashed lines. These quantities are defined by Eq. (3.26) in terms of respectively Eq. (3.25) and Eq. (3.24). The solid grey line shows the quantity $\delta_{{\rm SDK_{weak}}}^{2}/2$, which enters only in $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$.
@@
In the main panel we notice that the LO is dominated by the $WW$ “partonic” process, which has a much larger cross section than the $ZZ$ “partonic” process and therefore coincides with $2V$. Especially at 10 TeV, both $2V$ and $3V$ processes receive very large and negative contributions from NLO EW corrections. The coloured dotted lines in the main inset of the right plot of Fig. 23 indicates negative values for the lines depicted as solid, which we plot in absolute value in order to be visible on a logarithmic scale. It means that for both $2V$ and $3V$ processes, the relative NLO EW corrections are negative and larger than $100\%$ in absolute value. At 3 TeV they are smaller, but also in this case they are much larger than the contribution from double HBR production at LO, i.e. $4V$ at LO. Similarly to the case of the HBR, i.e. $3V$ at LO, which is smaller than the NLO EW corrections, the double HBR, i.e. $4V$ at LO, is smaller than the NLO EW corrections to $3V$ and so to the single HBR. Estimating the two-loop corrections, i.e. the NNLO to $2V$, as $\delta_{\rm NLO_{EW}}^{2}/2\simeq\delta_{{\rm SDK_{weak}}}^{2}/2$, we see that this contribution (the grey solid line) is even larger than the NLO EW corrections to the HBR contribution. In other words, looking at the case ${\cal O}(\alpha)$ and ${\cal O}(\alpha^{2})$, it appears that for a given $n$ the leading component of the ${\cal O}(\alpha^{n})$ corrections to LO in $2j_{\rm EW}$ production is the $n$-loop corrections, the subleading component is the (n1)(n-1)-loop corrections to HBR and so on; loops win over HBR.
@@ Appendix A ISR parton distribution
consistently with the renormalisation scheme that we will employ in the computations, with $G_{\mu}$ measured from the muon decay. Finally, as it is well known, the usage of LL-accurate ISR in NLO EW computations requires the inclusion, at the level of short-distance cross section, of additional terms required to attain formal NLO accuracy (treated on the same footage as change of factorisation-scheme contributions)  [^27] For more details, see e.g. Sec. 2 of Ref. Denner:2003iy or Appendix A of Ref. Bertone:2022ktl . . All our results at NLO will be computed including these terms, which are already available inside MadGraph5_aMC@NLO.